OpenStudy (anonymous):

determine a quadratic funtion with the given information; y-intercept of -4 and vertex (3,-7)

OpenStudy (ash2326):

A quadratic function or a parabola. Have you studied parabolas?

OpenStudy (eyust707):

Well a quadratic function is of the form: \[ax^2 + bx +c\]

OpenStudy (eyust707):

The y-intercept occurs when x is 0. Which means c must equal -4.

OpenStudy (eyust707):

The x coordinate of the vertex of a parabola is -b/2a. So: \[3={-b \over 2a}\] \[6a=-b\] \[-6a =b\] And the y coordinate is -7, So when we plug in a 3 we better get -7 \[f(3)= -7 = a(3)^2+b(3)-4 \] \[-7=a9+b3-4\] \[-3=9a+3b\] but before we said b must be -6a \[-3=9a+3(-6a)\] \[-3=9a-18a)\] \[-3=-9a\] \[a=1/3\] Then plug a back in to get b \[-6a =b\] \[-6({1 \over 3}) = b\] \[b=-2\] \[{1\over3}x^2 -2x -4\]

OpenStudy (anonymous):

thank you !!